[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Nov 09, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Nov 9 18:52:58 UTC 2016


HSU Shift Log for Wed Nov 09 2016 13:52:52 GMT-0500 (EST)

Shift: H

Forecaster: R Berg

Technical Issues Summary:

ATCF issues: A bunch of fixes from a previous EP92 are in the fix file for  
the new EP92.

N-AWIPS issues: The top right monitor on the Atlantic desk went out during  
the day. I notified Marc Laurent, and he was going to have Chris Mello look  
at it.

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161109/d8876709/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list